OpenStudy (anonymous):

find the quadratic poly.,the sum and the product of whose zeroes are 2 and -2/5 respectively

OpenStudy (anonymous):

\[ x^2 - S x + P =0 \]

OpenStudy (asnaseer):

if a quadratic has roots \(r_1\) and \(r_2\), then it can be written as:\[(x-r_1)(x-r_2)=0\]\[x^2-(r_1+r_2)x+r_1r_2=0\]

OpenStudy (asnaseer):

you should be able to solve from here

OpenStudy (anonymous):

oh god confusing

OpenStudy (asnaseer):

which part is confusing you?

OpenStudy (anonymous):

all :P

OpenStudy (asnaseer):

do you know what a quadratic equation is?

OpenStudy (anonymous):

\[ x^2 - 2 x - 2/5 = 0\\ 5 x^2 -10 x -2 =0 \]

OpenStudy (asnaseer):

If you don't then this site might help you understand: http://www.mathsisfun.com/algebra/quadratic-equation.html

OpenStudy (anonymous):

Also the sum and the products of the roots of \[ a x^2 + b x + c =0 \] are \[ S= - \frac ba\\ P= \frac c a \]

OpenStudy (anonymous):

i knw

OpenStudy (anonymous):

You can practice problems about roots of polynomials on my site http://www.saab.org/mathdrills/act.html

OpenStudy (aravindg):

am i late ? :P

OpenStudy (anonymous):

Did you understand my solution or @asnaseer which are the same

OpenStudy (aravindg):

@Parvathysubhash any further doubts on this?

OpenStudy (anonymous):

ysss @AravindG

OpenStudy (aravindg):

wat is it ?

OpenStudy (anonymous):

full can u show d entire step of dis question

OpenStudy (aravindg):

bt u see @eliassaab has done it very nicely here , can u tell me specifically where u hav doubt on?

OpenStudy (anonymous):

i knw its correct bt...??????????